OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

what's the equation of the circle? Remember it's (x-a)^2 + (y-b)^2 = r^2 right?

OpenStudy (anonymous):

Like in the previous question you just asked, \[(x-h)^2+(y-k)^2=r^2\] Where h is the x value (in this case 2), k is the y value (in this case -1), and r is the radius given (in this case 8). Plugging in those values, you get \[(x-2)^2+[y-(-1)]^2=8^2\] or \[(x-2)^2+(y+1)^2=64\]

OpenStudy (anonymous):

so what's the square of 8? 64 right? that eliminates 2 options already

OpenStudy (anonymous):

plug in values of a and b, which are 2 and -1

OpenStudy (anonymous):

x-2 and y-(-1)

OpenStudy (anonymous):

so must be answer B

OpenStudy (anonymous):

thanks tai,and everyone else.